please find the average value of the function:
f(x)=2x^2+3 on [0,2]


Answer by stanbon(48558)

please find the average value of the function:
f(x)=2x^2+3 on [0,2]
f(2)=2(4)+3=11
f(0)=3
Average = (11-3)/2=4
Cheers,
Stan H.